OpenStudy (anonymous):

4. What is the undefined value of 3b^2+13b+4/b+4? (1 point)

OpenStudy (anonymous):

It has a removable singularity at b=-4. \(3b^2+13b+4=(b+4)(3b+1)\) so \(\large \frac{3b^2+13b+4}{b+4}=3b+1\). Now plug in b=-4 and your undefined value is -11. I presume that's what you wanted.